OpenStudy (anonymous):

What is the fifth term of the sequence whose first term is 10 and whose common ratio is 1/2?

OpenStudy (anonymous):

Arithmetic or geometric sequence?

OpenStudy (anonymous):

i think geometric :/

OpenStudy (anonymous):

Do you multiply the first term by 1/2 to get the second term? If so, it is geometric. The equation to find the next term in the sequence is given by: (first term)*(Change)^(#term-1) so... \[10*(1/2)^{5-1}= 10*(1/2)^{4}\]

OpenStudy (anonymous):

so its 10 x 2 = 20?

OpenStudy (anonymous):

(1/2)^4= 1/16 and 1/16 * 10= 10/16 which simplifies to 5/8

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

np